OpenStudy (anonymous):

NO STRAIGHT ANSWERS PLZ help 2x + 5 = 27 A. 8.5 B. 16 C. 64 D. 11

pooja195 (pooja195):

subtract 5 from both sides

pooja195 (pooja195):

\[\huge~2x + 5-5 = 27-5\]

imqwerty (imqwerty):

sorry for the straight answers

OpenStudy (anonymous):

Simplifying 2x + 5 = 27 Reorder the terms: 5 + 2x = 27 Solving 5 + 2x = 27 Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '-5' to each side of the equation. 5 + -5 + 2x = 27 + -5 Combine like terms: 5 + -5 = 0 0 + 2x = 27 + -5 2x = 27 + -5 Combine like terms: 27 + -5 = 22 2x = 22 Divide each side by '2'

OpenStudy (anonymous):

its ok i got a warning from the moderator

OpenStudy (anonymous):

first you want to do x/2 for both sides then subtract 5 from both sides and you have your answer

OpenStudy (anonymous):

ok thank you @CatGamer02

OpenStudy (anonymous):

no prob :)